Dragsters can achieve an average acceleration of 26.0 m/s2. Suppose a dragster accelerates from rest at this rate for 6s. How far does it travel in this time?


1
Expert's answer
2021-08-02T08:42:13-0400
"d=v_0t+\dfrac{1}{2}at^2,""d=0+\dfrac{1}{2}\cdot26\ \dfrac{m}{s^2}\cdot(6\ s)^2=468\ m."